### v3

# Adaptive In-conversation Team Building for Language Model Agents

Abstract
Leveraging multiple large language model (LLM) agents has shown to be a promising approach for tackling complex tasks, while the effective design of multiple agents for a particular application remains an art. It is thus intriguing to answer a critical question: Given a task, how can we build a team of LLM agents to solve it effectively? Our new adaptive team-building paradigm offers a flexible solution, realized through a novel agent design named Captain Agent. It dynamically forms and manages teams for each step of a task-solving process, utilizing nested group conversations and reflection to ensure diverse expertise and prevent stereotypical outputs, allowing for a flexible yet structured approach to problem-solving. A comprehensive evaluation across six real-world scenarios demonstrates that Captain Agent significantly outperforms existing multi-agent methods with 21.94% improvement in average accuracy, providing outstanding performance without requiring task-specific prompt engineering. Our exploration of different backbone LLM and cost analysis further shows that Captain Agent can improve the conversation quality of weak LLM and achieve competitive performance with extremely low cost, which illuminates the application of multi-agent systems.
[…]
1 Introduction
The success of large language model (LLM) agents (Yao et al., 2022; Yang et al., 2023a; Furuta et al., 2024; Yang et al., 2024a; Hong et al., 2024) with its outstanding in-context learning (Dong et al., 2022; Brown et al., 2020; Yang et al., 2023b; Dai et al., 2023; Li et al., 2023c), planning (Sun et al., 2024; Xie et al., 2024; Liu et al., 2023a; Valmeekam et al., 2022; Wei et al., 2022a; Yuan et al., 2023b; Zheng et al., 2024), tool-using (Qin et al., 2023a, b; Schick et al., 2024; Cai et al., 2023; Yuan et al., 2023a; Paranjape et al., 2023; Zhang et al., 2024b; Huang et al., 2023; Ma et al., 2024), and conversation (Fernandes et al., 2023; Wang et al., 2023c; Yang et al., 2024b) capabilities allow us to relate human’s team building and collaboration abilities to the multiple language model agents (multi-agent) system (Wang et al., 2023a; Xi et al., 2023; Wu et al., 2023; Suzgun & Kalai, 2024a; Hong et al., 2023; Zhang et al., 2024b, 2023a; Valmeekam et al., 2023; Wang et al., 2024; Saha et al., 2023; Liang et al., 2023; Du et al., 2023; Chen et al., 2024). Humans have developed abilities that enable us to form teams and effectively solve problems. These abilities are rooted in communication, social cognition, problem-solving and decision-making, social learning and imitation, and shared intentionality (Elimari & Lafargue, 2020; Confer et al., 2010). The interplay of the above abilities allows people to organize different teams for problems to ensure that tasks are completed successfully, which brings us to a critical question in a multi-agent system:
Given a task, how can we build a team of LLM agents to solve it effectively?
A straightforward paradigm would be to build a static agent team beforehand based on the task instruction and let them solve the task collaboratively (Chen et al., 2023; Wu et al., 2023). However, static build necessitates maintaining a team with all the required expertise for the whole task cycle. As the complexity of the task increases, the total number of team members may grow significantly. Always proceeding with such a large team makes it challenging to manage the team members effectively and efficiently. Furthermore, static teams may lack the adaptability to respond to dynamic changes in task requirements or unforeseen challenges. Imagine a prehistoric human tribe: was everyone involved in every task? The answer is unlikely affirmative. Those responsible for hunting may not participate in medical care and those responsible for cooking may not involve themselves in management. The major task, survival, was ensured by each individual group sticking to their roles and subtasks. In fact, when human organizations handle a complex task, we tend to form multiple teams for each subtask at different stages of the task-solving procedure, which still guarantees a diverse set of expertise is leveraged demanded by the task complexity (Mao et al., 2016).
Inspired by how humans assemble teams for a complex task, we introduce a new multi-agent team-building paradigm: adaptive build. This paradigm facilitates the flexible assembly of agents with specific skills and knowledge as demands evolve in the process of task-solving. To realize this paradigm, we propose a new adaptive builder agent, Captain Agent, to build, manage, and maintain agent teams for each problem-solving step in the conversation. Captain Agent has two core components: (1) adaptive multi-agent team building and (2) nested group conversation and reflection. Captain Agent will communicate with a User Proxy, who can provide the general task instructions at the beginning. When assigned a task, Captain Agent begins by formulating a strategic plan. This plan involves a cyclical process that continues until the task is successfully completed. In the first phase of the cycle, Captain Agent identifies a specific subtask, outlines the necessary roles, and assembles a team of agents equipped with the appropriate tools. In the subsequent phase, this team engages in a dialogue with a versatile tool to address the subtask. Upon completion, a reflector LLM reviews the process and provides Captain Agent with a detailed reflection report. Based on this feedback, Captain Agent either adjusts the team composition or the subtask instructions and repeats the cycle or concludes the task and presents the final outcomes.
We evaluate state-of-the-art multi-agent approaches for complex task solving and our adaptive build approach with Captain Agent on six real-world scenarios, including many mathematics problem-solving (Hendrycks et al., 2021b), data analysis (Hu et al., 2024b), programming (Le et al., 2020), scientific problem-solving (Wang et al., 2023b) (Physics and Chemistry), and information retrieval (Mialon et al., 2024). Our experimental results demonstrated the outstanding ability of Captain Agent in various scenarios without prompt engineering. Captain Agent achieves distinguishing results compared to other single and multi-agent methods and frameworks when using the same prompt for each task, with an average of 21.94% improvement on average accuracy. Ablation studies on static and adaptive building paradigms show that the adaptive team outperforms the static team in four of five scenarios (and matches in one scenario), exhibiting the superiority of the adaptive build paradigm across different scenarios. We also demonstrated that handcraft agents and handcraft tools contribute equally to the final results. We further explore the influence of different backbone LLM for both Captain Agent and nested group chat members or only for nested group chat members. We observe that: (1) Captain Agent with a strong backbone can improve the quality of nested group chat in which the members equipped with weak backbone, and (2) a small model with distinguishable instruction following ability can achieve outstanding performance with low cost.
2 Captain Agent
The overall workflow of Captain Agent is illustrated in Figure 1. We first establish a main conversation between the user and Captain Agent. When receiving a user task, Captain Agent will: (Step 1) first decompose the task into a list of subtasks and suggest several roles for this subtask. A team of agents will be created accordingly. Each agent will be equipped with several predefined tools according to the abilities defined by Captain Agent (Section 2.1); (Step 2) the agent team will attempt to solve a subtask via a nested conversation. Once the nested conversation reaches the terminal condition, the conversation history will be summarized and analyzed by a reflector LLM, and the final result will be feedback to Captain Agent (Section 2.2). These two steps will repeat certain rounds in the main conversation until reaching the max round limitation or terminal condition.
2.1 Adaptive Multi-agent Team Building
After identifying a subtask in Step 1 following a corresponding prompt, Captain Agent will list several roles for the subtask. These roles will then pass into a retrieval, selection, and generation process guided by Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Ram et al., 2023). Created agents will be equipped with a well-designed profile (system message111System message refers to a profile that defines an agent’s persona and task-specific instructions.) and high-quality tools.
Agent and tool retrieval. Captain Agent will suggest required roles for the incoming user task at round . includes a short sentence that can describe itself. We then retrieve top- agents and top- tools for according to the embedding similarity between and the agent/tool description in the library, as follows:
where and are the amount of retrieved agents and tools from agent library and tool library , respectively. denotes the extraction of the embeddings. We use to denote the word ”candidate.” We then equip with corresponding .
Agent filtering. Finding an exactly matched agent from our agent library for every real-world task is nearly impossible. Therefore, we prompt an LLM-based filter to select the top- suitable agent for from , simulating the human selection process from massive retrieved results (e.g., searching on the internet). We provide candidate agents , suggested role , and the user task as input to the filter LLM. The filter LLM is required to return the most suitable one agent from . An abstention mechanism is adopted to this process, i.e., the filter LLM can return None to prevent irrelevant agents from being selected as team members. This often happens when is small. We then go through an agent generation process for those with no suitable agents.
Agent generation. We design a step-by-step and templated agent generation process for with no suitable agent selected at the previous step, motivated by Wei et al. (2022b); Shinn et al. (2024). Generally, it includes three steps: (1) generate a basic profile, including name, persona, and required skills as the system message; (2) combine the basic system message with predefined general task-solving and environment interaction instructions, for example, group chat information; (3) retrieved suitable tools from according to the generated system message and bind the new agent with retrieved tools. We then update the newly generated agent into for future use.
Team memory. Captain Agent will cache the built team into its conversation-wise temporary memory, which only includes basic details for each member of the team and no previous conversation history. Captain Agent can prompt the same team with new instructions to seek an alternate result without going through the team-building process, which helps reduce the total cost.
2.2 Nested Conversation and Reflection
Once the team-building process is completed, a new conversation will be established between the created agent team members apart from the main conversation, which we call ”nested conversation.” Captain Agent will provide a subtask to the nested conversation. At the end of the nested conversation, an LLM-based reflector will analyze the whole conversation history. Captain Agent next-step decision (continue or conclude) will based on the reflector’s report.
Nested conversation. In the nested conversation, an LLM-based conversation manager will adaptively select the next speaker from the agent team members based on the conversation history. Team members will share the same conversation history in the nested group chat, but they cannot access the history from the main conversation. Conversation history consists of the previous agent’s action and corresponding feedback. Each agent’s action can be natural language instruction and/or a Python code block. Specifically, we use raw Python function as our tool format, which allows agents to create new tools with their assigned tools by incorporating into a self-written Python code block, which helps extend agent’s action set creatively.
Conversation reflection. A conversation can easily escalate into a conflict once an agent starts a dispute with another, where they will never make a consensus and continue until reaching the max-round. This often happens when Captain Agent provides an ambiguous instruction or the subtask is not clear enough. Therefore, we design a reflection mechanism for nested conversation. An LLM-based reflector will summarize the conversation history and check all disputes inside the conversation, then provide feedback to the Captain Agent with a reflection report. Captain Agent will make further decisions based on this report. All reflection reports will be recorded in the main conversation.
[…]
| Mathematics problems | MATH (Hendrycks et al., 2021a) | 196 |
| Programming | HumanEval (Chen et al., 2021) | 164 |
| Data Analysis | DABench (Hu et al., 2024a) | 257 |
| information retrieval | GAIA (Mialon et al., 2023) | 165 |
| (Scientific) Chemistry | SciBench (Wang et al., 2023b) | 41 |
| (Scientific) Physics | SciBench (Wang et al., 2023b) | 34 |
2.3 Benefits over Static Team
A static team may limit the team’s ability coverage. Although building a large number of agents with comprehensive persona or skill sets can address the limitation in ability coverage, it is challenging for LLMs to handle a long context that introduces all the participant members. Unexpectedly long contexts will primarily reduce the quality of the conversation. Meanwhile, agents with redundant functionality will also be involved in the task-solving process. In contrast, Captain Agent can adaptively select and build more optimized agent teams for the current task, reducing the prompting load for LLMs and redundant output from irrelevant agents without sacrificing the diversity in the team.
3 Evaluation
3.1 Experimental Setup
Scenarios and datasets. We choose six real-world scenarios, including mathematics, programming, data analysis, information retrieval, and science problem-solving. Each scenario was chosen for its unique ability to demonstrate specific capabilities and performance metrics of the agent systems. This ensures a holistic assessment of Captain Agent against the baselines across various critical dimensions of computational and cognitive skills. We bind each scenario with a challenging open-source dataset, as shown in Table 1. Due to cost limitations, we sample a subset of MATH according to its original distribution of each question type.
Compared methods and implementation. For mathematics problems, programming, data analysis, and scientific scenarios, we investigate the performance of Captain Agent and four different methods, including Vanilla LLM (prompt an LLM once for an answer), AutoAgents (Chen et al., 2023), Meta-prompting (Suzgun & Kalai, 2024a), AgentVerse (Chen et al., 2024), DyLAN (Liu et al., 2023b), and a two-agent system (a system involving an Assistant agent with an Executor agent) realized with AutoGen (Wu et al., 2023). Specifically, we implement AutoAgents with AutoGen as the official implementation is unstable and unsuitable for large-scale experiments. For meta-prompting, we improve its code execution ability by reproducing a real-time Python interpreter. All these methods will be prompted with the same task-specific instruction (refer to Appendix E).
For information retrieval scenarios, we compare Captain Agent with the top-5 baselines (with reference) reported to the GAIA validation leaderboard, which includes AutoGen: GAIA_Orchestrator (a specific three-agent setting organized by an Orchestrator agent designed for GAIA) (GAIA_Orchestrator, 2024), FRIDAY (Wu et al., 2024), Warm-up Act222Warm-up Act has no open-sourced implementation., GAIA official implementation, and AutoGPT (Yang et al., 2023a).
For Captain Agent, we adopt all-mpnet-base-v2 to calculate the sentence embedding for agent and tool retrieval. Our environment will communicate with Captain Agent by providing the feedback of code execution, tool calling (adaptive build), nested conversation reflection results, and a default reply: I’m a proxy, and I can only execute your code and tool or end the conversation. If you think the problem is solved, please reply to me only with ’TERMINATE.’
| Method | Mathematics | Programming | Data Analysis | (Sci) Chemistry | (Sci) Physics | Avg. |
| Vanilla LLM | 51.53 | 84.76 | 6.61 | 39.02 | 31.25 | 40.98 |
| Meta-prompting | 68.88 | 19.51 | 39.69 | 41.46 | 43.75 | 43.47 |
| AutoAgents | 56.12 | 84.76 | 57.98 | 60.98 | 50.00 | 63.58 |
| DyLAN | 62.24 | 90.24 | - | 45.45 | 51.16 | - |
| AgentVerse | 69.38 | 42.68 | - | 42.42 | 37.21 | - |
| AutoGen: Assistant + Executor | 74.49 | 93.90 | 82.88 | 60.98 | 43.75 | 79.89 |
| Captain Agent | 77.55 | 96.95 | 88.32 | 65.85 | 53.12 | 84.25 |
| Method | Level 1 | Level 2 | Level 3 | Avg. |
| AutoGPT (gpt-4) | 13.21 | 0.00 | 3.85 | 4.85 |
| gpt-4-turbo | 20.75 | 5.81 | 0.00 | 9.70 |
| gpt-4-turbo + manually selected plugins | 30.30 | 9.70 | 0.00 | 14.6 |
| Warm-up Act (gpt-4-turbo) | 35.19 | 15.12 | 0 | 17.58 |
| FRIDAY (gpt-4-turbo) | 45.28 | 34.88 | 11.54 | 34.55 |
| AutoGen: GAIA_Orchestrator (gpt-4-turbo) | 54.72 | 38.31 | 11.54 | 39.39 |
| Captain Agent (gpt-4-turbo) | 56.60 | 39.53 | 11.54 | 40.60 |
Agent and tool library. We initialize our agent library based on a small subset of problem instances from each dataset (described in Section 3.4, about 20 questions for each dataset) in Table 1. Specifically, we run Captain Agent on the subset and iteratively update the library by adding the generated agents and keeping our agent library unchanged during the main experiment. Our agent library also supports all hand-crafted agents (of the ConversableAgent class) archived in AutoGen (details in Appendix F). All these agents follow the ConversableAgent interface to converse with each other. Our tool library consists of a suite of callable Python functions intended for freeform coding. The agents can freely import functions from the tool library and write free-form code to integrate the outputs to handle sophisticated tasks (see also Appendix C and G). The library contains three main categories of tools: math, data analysis, and information retrieval. For each category, we summarize the patterns of the corresponding dataset and manually craft a set of functions that suit the tasks.
3.2 Evaluation Protocol
For mathematics, data analysis, and science scenarios, we report the accuracy of each method by comparing the final result from each method and ground truth. To ensure fairness in evaluation, we transform different result formats into a uniform format, preventing the correct answer from being judged incorrect due to format mismatches. For programming scenarios, we run the code provided from each method and output a unique token if the code successfully passes all tests. We then count the success token and calculate the accuracy for each method.
3.3 Main Results
Table 2 and 3 report the comparison results between Captain Agent and eight different baselines on six real-world scenarios. Baseline results on information retrieval are extracted directly from the GAIA leaderboard.
Findings 1: Diverse agents can help trigger accurate expertise output for problem-solving. By comparing the results from Captain Agent, AutoAgents, and AutoGen Assistant + Executor, we observe that Captain Agent and AutoAgents averagely outperform AutoGen Assistant + Executor on (Sci) Chemistry and (Sci) Physics scenarios. These scenarios required expertise knowledge, which the AutoGen Assistant with a fixed system message is hard to complete. Captain Agent and AutoAgents can create diverse experts by assigning different domain-specific system messages to agents, which helps better trigger the intrinsic knowledge inside an LLM to provide an accurate answer. Captain Agent outperforms AutoAgents in all the scenarios because Captain Agent can provide a high-level plan and solve each step with adaptive instructions and an agent team.
Findings 2: Adaptive team-building boosts performance with no task preference. It is obvious that Captain Agent achieves outstanding results over all scenarios, indicating that Captain Agent is free from task preference. Incorporating different agents into the team at a proper time gives Captain Agent the ability to solve difficult tasks like science and information retrieval problems step-by-step. We observe that Meta-prompting fails in science scenarios due to the inability to decompose science problems into the fine-grain subtasks that one agent can solve. Captain Agent with the agent-team building paradigm neither requires a task that can be decomposed into a subtask that can only be solved by an agent nor requires all agents to be involved in the conversation. Compared with other agent-team building methods such as AutoAgents, DyLAN, and AgentVerse, Captain Agent can not only We further discuss the static and adaptive teams in Section 3.4.1.
3.4 Analysis and Ablation Studies
In this section, we dive into the difference between static and adaptive team-building, the influence of agent and tool libraries, and the possibility of working with open-weight models. We perform ablation studies on a subset from Table 1. Specifically, we choose 17 problems from MATH and 25 problems from HumanEval according to the AutoGenBench (AutoGenBench, 2024), in which the problems are randomly selected from GPT-4 failure set. For DABench, we randomly selected 25 problems, and for SciBench, we randomly selected 19 problems for chemistry and physics according to the number of textbooks. The evaluation protocol is the same as in Section 3.3.
3.4.1 Static vs. adaptive team-building
To further explore how building a team adaptively in conversation impacts the overall performance, we compare Captain Agent in two settings: (1) building a team at round 1 in the main conversation without changing the team members (static team) and (2) adaptively changing team members and instructions across all rounds in the main conversation (adaptive team). The comparison results in Figure 2 show that except in (Sci) Physics scenario, adaptive build outperforms static build across other scenarios. On data analysis scenarios, adaptive build outperforms static build by over 30.43% on accuracy, showing the importance of changing team members when dealing with data analysis scenarios.
| Captain Agent | Information Retrieval | ||||
| Agent Library | Tool Library | Level 1 | Level 2 | Level 3 | Avg. |
| - | - | 32.07 | 13.95 | 3.84 | 18.18 |
| ✓ | - | 37.73 | 30.23 | 7.69 | 29.09 |
| - | ✓ | 39.62 | 19.78 | 7.69 | 24.24 |
| ✓ | ✓ | 56.60 | 39.53 | 11.54 | 40.60 |
3.4.2 Tool library and agent library
To demonstrate the effectiveness of agent and tool library, we conduct ablation studies over three settings: (1) remove both agent and tool library, (2) remove agent library while retaining the tool library, (3) remove tool library while retaining agent library, and (4) keep both agent and tool libraries. We selected GAIA as our target dataset because it represents the strong requirement for high-quality tools and the diversity of problem-solving abilities in different domains, which perfectly match the purpose of this ablation study. As shown in Table 4, removing the agent library and tool library can both significantly impair the system’s performance. While both the tool and agent libraries can enhance performance independently, optimal results are achieved only when both libraries are employed concurrently. Handling level 1 tasks requires a moderate amount of web browsing and reasoning steps, which can be achieved by several single-turn tool calls or experts writing and executing code iteratively. Notably, without an agent library, Captain Agent performs much worse on Level 2 tasks. This is because these tasks are more sophisticated and mostly involve a significant number of web navigation and reasoning steps. Web browsing involves complex and dynamic interactions that are poorly suited to static tool libraries. The tasks require agents to coordinate multiple tools to solve them, which is a process prone to error in web scenarios filled with uncertainty.
3.4.3 Ablation on Reflector
Reflection is one of the most important components in Captain Agent as its feedback represents the essence of the whole nested conversation and will, therefore, largely impact the decision-making process of Captain Agent. In Figure 3, we statistic the performance of Captain Agent with and without reflection and the success rate of the reflector. We define the success rate as: if the LLM-based reflector can detect the dispute in nested conversation and highlight it for Captain Agent to avoid potential next-step error, we will mark it as ”success.” We calculate the success rate of the reflector across all problems under the corresponding scenario. We observe that in (sci) physics problems, the reflector represented a vital role with a high success rate and helped improve the performance of Captain Agent, but the performance of Captain Agent under physics remains lowest across all other scenarios. This is because the reflector cannot refine the problem-solving process but can detect the conflict inside a nested conversation and infer the possible error that may be incurred by the conflict. Therefore, the reflection mechanism helps prevent error propagation in the multi-round problem-solving process designed for Captain Agent.
3.4.4 Ablation on LLM Backbone and Cost Analysis
The high token cost associated with LLMs has always been a significant barrier to the practical deployment of agents, rendering them economically unfeasible. However, with the development of model distillation, the small model can also achieve competitive results with an agent framework on complex tasks. To evaluate Captain Agent’s ability to generalize to different LLM backbones for cost reduction, we equip Captain Agent and its nested experts with four different backbones, including gpt-4-0125-preview, gpt-4o-mini, LLaMA-3-70B-Instruct, and LLaMA-3-8B-Instruct, and compare it with all the baselines equipped with gpt-4-0125-preview. Our cost statistic covers the whole Captain Agent workflow, including Captain Agent output, performing agent and tool selection, agent generation, and nested group conversation. We summarize the results based on the ranking of each setting in Figure 4. We observe that (1) Captain Agent with the same backbone as baselines can achieve the best performance but also the highest cost; (2) Captain Agent with gpt-4o-mini outperforms all other baselines but with a significantly lower cost, revealing the possibility of using a small language model as the backbone of an agent framework without seeing cost as a major issue.
3.4.5 Agent Selected in Each Scenario
Our agent library records 541 agents in total after we complete all experiments, and we summarized the top-10 selected agent’s frequency for each scenario in Figure 5. It is obvious that selected agents are highly related to the corresponding scenarios because we prompt Captain Agent in the system message to create a verification role to keep the conversation going smoothly. We also notice that in some scenarios (mathematics, data analysis, and programming), agents with a non-specific name and description will have a high selection rate (e.g., PythonMath_Expert, DataAnalysis_Expert, CodeReview_Expert, etc.), as the embedding similarity between their description and user task is generally high. But we can still observe some specific agents being selected, like AppliedMath_Expert, StringManipulation_Expert, and NumPySciPy_Expert, etc. For the information retrieval scenario, WebResearch_Expert has an extremely high selection rate. This agent is one of our hand-crafted agents who can search the internet and summarize a page in Markdown format, and Agent Filter can select this agent with Captain Agent instruction properly.
4 Related Work
Large language models (LLMs) represent a significant advancement in artificial intelligence, showcasing remarkable capabilities in various aspects, including reasoning (Wei et al., 2022b; Yao et al., 2024; Morishita et al., 2023; Zhang et al., 2023b; Li et al., 2023a; Ho et al., 2022), planning (BabyAGI, 2023; Song et al., 2023; Valmeekam et al., 2023; Liu et al., 2023b), and adaptability to novel real-world observations (Shi et al., 2024; Hong et al., 2023; Yang et al., 2023a; Dan et al., 2023; Zhou et al., 2023a; Bharadhwaj et al., 2023). Leveraging the inherent versatility of LLMs as generalized models adaptable to diverse scenarios, numerous efforts have been dedicated to the development of intelligent agents (Wu et al., 2023; Xi et al., 2023; Zhang et al., 2024b; Sumers et al., 2023; Zhou et al., 2023b) where LLMs serve as foundational components. For instance, one typical algorithm, React (Yao et al., 2022), employs one single LLM to iteratively generate both reasoning trajectories and task-specific actions. This interleaved process enables the agent to engage in dynamic reasoning. In addition, LLM agents can also harness external tools (Qin et al., 2023a, b; Schick et al., 2024; Cai et al., 2023; Yuan et al., 2023a; Paranjape et al., 2023; Zhang et al., 2024b; Huang et al., 2023; Ma et al., 2024), leveraging both their internal capabilities and external resources, collaborating effectively to solve more intricate problems.
The success of a single-agent system motivates the development of multiple-agent systems (Wang et al., 2023a; Xi et al., 2023; Chen et al., 2023; Wu et al., 2023; Suzgun & Kalai, 2024a; Hong et al., 2023; Zhang et al., 2024b, 2023a; Valmeekam et al., 2023; Wang et al., 2024; Saha et al., 2023; Liang et al., 2023; Du et al., 2023). Methods focusing on static build require a protocol for agents to communicate with each other in a group chat and a builder that can receive the user’s instruction and output an agent list (Wu et al., 2023; Chen et al., 2023; Hong et al., 2023). The builder can be a human (Wu et al., 2023; Hong et al., 2023) or a LLM agent (Chen et al., 2023). There are other works breaking down complex tasks into smaller components, each of which is then handled by a single specialized agent with detailed natural-language instructions (Suzgun & Kalai, 2024b; Zhuge et al., 2023). This task decomposition reduces the prediction burden on each agent by avoiding irrelevant context. For instance, meta-prompting (Suzgun & Kalai, 2024b) involves a meta-model decomposing tasks and assigning subtasks to different LLMs for completion and aggregation. However, these methods lack a backtracking mechanism, making the problem-solving prototype fixed once agents are generated.
5 Conclusion and Discussion
We introduce a new paradigm for multi-agent team-building, adaptive build. This new paradigm helps ensure diversity and prevent limited knowledge extraction. The new paradigm executed by our proposed agent, Captain Agent, manages agent teams for problem-solving steps using adaptive multi-agent team building and nested group conversation and reflection. Experimental results across six real-world scenarios demonstrate Captain Agent’s efficacy in various tasks without prompt engineering, achieving superior results compared to existing methods. Ablation studies confirm that each component contributes equally to overall performance, underscoring the robustness of our approach.
Impact Statement
Our method dynamically ensembles LLM agents and equips them with versatile tools, allowing them to efficiently and effectively solve complex tasks. However, the development of agent systems that interact with the web environment raises safety concerns. The scope of our experiment in real-world interaction is limited to solving GAIA tasks, where the agents are required to search the web and browse websites. The agents are restricted from accessing publicly available information and are not capable of publishing content on the web. This ensures that our experiments remain both non-invasive and safe.
Appendix A Limitations
The first limitation of our work is cost. A conversation involving the GPT-4 model costs more than a single-agent system. Although we have reduced the cost by decreasing the participant nested group chat agents, it still has a large conversation and profile as context input. The trade-off between performance and cost will become one of the possible future works for further exploration, like window context, conversation pruning, or conversation compression. Another limitation of our work is the lack of thinking about model selection. We should also notice that the current evaluation of LLM is not perfect. Data leaking is widespread in the pertaining process and will cause the misalignment between the test and real-world performance (Zhang et al., 2024a; Xu et al., 2024). Therefore, a comprehensive yet fair evaluation is important for us to further discuss the ability of model diversity. In this work, all agents are communicated under AutoGen with a highly customized communication framework. This limited us to generalize Captain Agent to other existing hand-crafted agents (i.e., absorb those fantastic agents into our agent library). Therefore, a universal agent communication protocol is important.
Appendix B Difference Between Existing Team-building Frameworks
| Math | Prog. | DA | (Sci) Phys. | (Sci) Chem. | Avg. Rank | ||
| Backbone | Performance (Accuracy, higher is better) | ||||||
| Vanilla LLM (gpt-4-0125-preview) | 52.94 | 72.00 | - | 26.32 | 31.58 | 6.8 | |
[…]
| AutoAgent (gpt-4-0125-preview) | 64.71 | 88.00 | 52.17 | 47.37 | 68.42 | 3.2 | |
[…]
| AgentVerse (gpt-4-0125-preview) | 64.71 | 20.00 | - | 36.84 | 42.11 | - | |
[…]
| w/ Llama-3-70B-Instruct | 47.06 | 80.00 | 56.52 | 43.75 | 36.84 | 4.6 | |
| Captain Agent | w/ Llama-3-8B-Instruct | 5.89 | 48.00 | 34.78 | 5.26 | 5.26 | 7.4 |
| Backbone | Cost for Task Completion (US Dollars, lower is better) | ||||||
[…]
| Two-Agents (gpt-4-0125-preview) | 3.10 | 2.82 | 5.32 | 1.34 | 2.33 | 5.2 | |
| Meta-prompting (gpt-4-0125-preview) | 2.92 | 9.88 | 8.64 | 4.18 | 4.96 | 5.8 | |
| AutoAgent (gpt-4-0125-preview) | 4.59 | 18.32 | 33.58 | 12.48 | 12.28 | 7 | |
| DyLAN (gpt-4-0125-preview) | 3.01 | 8.76 | - | 7.10 | 8.07 | - | |
| AgentVerse (gpt-4-0125-preview) | 7.63 | 13.59 | - | 26.34 | 23.56 | - | |
| w/ gpt-4-0125-preview | 7.95 | 23.67 | 39.88 | 15.21 | 18.68 | 8 | |
[…]
| w/ Llama-3-70B-Instruct | 0.89 | 1.92 | 0.89 | 1.18 | 1.48 | 3.4 | |
| Captain Agent | w/ Llama-3-8B-Instruct | 0.05 | 0.03 | 0.02 | 0.06 | 0.08 | 1 |
In this section, we will discuss the difference between Captain Agent and other famous agent team-building frameworks, including AutoAgent (Chen et al., 2023) AgentVerse (Chen et al., 2024), and DyLAN (Liu et al., 2023b).
Difference between AgentVerse and Captain Agent
Compared with Agentverse, Captain Agent supports more flexible agent team building and collaboration. AgentVerse includes two types of framework: dynamic team and handcrafted team. The dynamic team completes part of the tasks with the recruitment process, in which some agents are recruited in a fixed process (recruit – chat or comment – evaluate – reflect), and the handcrafted team completes other tasks without the recruitment process. In contrast, we did not design fixed teams for any tasks. Moreover, unlike the fixed sequential process, Captain Agent can also be involved in the nested group chat as it can solve part of the problems by itself and pass the solution into the nested chat. Furthermore, the Captain Agent can cache teams in its memory and call back at a proper time. Therefore, the Captain Agent acts like a time leaper who can participate in different teams on different timelines to help derive better solutions.
Difference between DyLAN and Captain Agent
DyLAN is a static build process in which the multi-agent debate starts with a fixed and manually predefined group of experts. On the other hand, DyLAN handcrafts a pool of expert names, their corresponding prompts, and their capabilities. The agent selection from pool to expert group member is manually performed. The framework requires manual effort to function properly.
Appendix C Case Study Comparing Captain Agent and Others
We demonstrate the benefits and functionalities of Captain Agent in two distinct scenarios. The first scenario involves the automatic tailoring of specialized agentic systems for solving complex user queries. The second scenario highlights its ability to replace manually designed agent systems without human intervention, achieving similar performance to the handcrafted three-agent system OptiGuide (Li et al., 2023b) in a specific supply chain optimization scenario.
C.1 Case Study 1: Arbitrary User Question Answering
As illustrated in Figure 6, a user seeks specific information from an article by Emily Midkiff, published in June 2014. In response, Captain Agent assembles a team of specialized agents: WebBrowsing_Expert, Myth_Expert, and Verification_Expert. Each agent is designed to address a specific domain, and together they effectively collaborate to solve the problem, under the coordination of Captain Agent. In contrast, a simple two-agent system from AutoGen fails to resolve the issue due to limitations in the Assistant’s capabilities. This highlights Captain Agent’s ability to build an optimal team of agents and ensure their collaboration, whereas basic two-agent systems may lack the necessary functionalities to address more complex queries.
C.2 Case Study 2: Domain Specific Question Answering
Figure 7 illustrates a comparison between Captain Agent and OptiGuide (Li et al., 2023b), a handcrafted three-agent system designed for supply chain optimization. When provided with the necessary background information, Captain Agent is able to generate a similar group structure, successfully solving the task. In contrast, an examination of the implementation code reveals that OptiGuide includes a fixed structure requiring over 100 lines of code, meticulously crafted by humans. In comparison, CaptainAgent accomplishes the same task with only a few lines of code, significantly reducing the complexity for developers and easing domain-specific development.
Appendix D Instruction of Captain Agent
We design a general profile message (system message) for Captain Agent to ensure it can execute our paradigm efficiently and effectively. Instructions are in markdown format, including a planning instruction that can decompose the task into multiple steps, a building instruction (the seek_experts_help), a post-seek_agent_help instruction, and some general instructions that help task solving.
D.1 System Message
D.2 Reflector LLM
D.3 Agent Selector LLM
Appendix E Task Instructions
We design instructions manually for each scenario and ensure all baselines and Captain Agent receive the same instructions for a fair comparison333Except for the information retrieval scenario (GAIA), in which we use the results directly from the leaderboard.. All instructions include the basic information of the scenario and may suggest some possible Python libraries, including pandas, numpy, scipy, and sympy.
E.1 Instruction for Mathematics
E.2 Instruction for Programming
E.3 Instruction for Data Analysis
E.4 Instruction for Science (Chemistry and Physics)
E.5 Instruction for World-information Retreival
Appendix F Agent Library
We provide an example of the agent recorded in the agent library below:
Appendix G Tool Library
This section provides the names and descriptions of our manually created tool library. The tools are categorized into three classes: Information Retrieval, Data Analysis and Math Problem Solving. For each category, we summarize the patterns of the corresponding dataset and manually craft a set of functions suits the tasks and can potentially enhance the agents’ task resolution capability.
| Tools | Description |
| scrape_wikipedia_tables | Scrapes Wikipedia tables based on a given URL and header keyword. |
| transcribe_audio_file | Transcribes the audio file located at the given file path. |
| youtube_download | Downloads a YouTube video and returns the download link. |
| academic_search | Perform an academic search of papers, authors or an author’s papers. |
| docx_to_md | Converts a DOCX file to Markdown format. |
| pptx_to_md | Convert a PowerPoint presentation (PPTX) to Markdown format. |
| spreadsheet_to_md | Convert an Excel spreadsheet file to Markdown format. |
| extract_pdf_image | Extracts images from a PDF file and saves them to the specified output directory. |
| extract_pdf_text | Extracts text from a specified page or the entire PDF file. |
| get_youtube_caption | Retrieves the captions for a YouTube video. |
| image_qa | Answers your questions about a given image. |
| optical_character_recognition | Perform optical character recognition (OCR) on the given image. |
| perform_web_question_answering | Perform web search according to keyword and answer your question on each webpage search result, or directly on the webpage if the keyword is a URL. For each search result, a response to the question is provided. |
| scrape_wikipedia_tables | Scrapes Wikipedia tables based on a given URL and header keyword. |
| Tools | Description |
| calculate_correlation | Calculate the correlation between two columns in a CSV file. |
| calculate_skewness_and_kurtosis | Calculate the skewness and kurtosis of a specified column in a CSV file. The kurtosis is calculated using the Fisher definition. |
| detect_outlier_iqr | Detect outliers in a specified column of a CSV file using the IQR method. |
| detect_outlier_zscore | Detect outliers in a CSV file based on a specified column. The outliers are determined by calculating the z-score of the data points in the column. |
| explore_csv | Reads a CSV file and prints the column names, shape, data types, and the first few lines of data. |
| shapiro_wilk_test | Perform the Shapiro-Wilk test on a specified column of a CSV file. |
| Tools | Description |
| calculate_circle_area_from_diameter | Calculate the area of a circle given its diameter. |
| calculate_day_of_the_week | Calculates the day of the week after a given number of days starting from a specified day. |
| calculate_fraction_sum | Calculates the sum of two fractions and returns the result as a mixed number. |
| calculate_matrix_power | Calculate the power of a given matrix. |
| calculate_reflected_point | Calculates the reflection point of a given point about the line y=x. |
| complex_numbers_product | Calculates the product of a list of complex numbers. |
| compute_currency_conversion | Compute the currency conversion of the given amount using the provided exchange rate. |
| count_distinct_permutations | Counts the number of distinct permutations of a sequence where items may be indistinguishable. |
| evaluate_expression | Evaluates a mathematical expression with support for floor function notation and power notation. |
| find_continuity_point | Find the value that ensures the continuity of a piecewise function at a given point. |
| fraction_to_mixed_numbers | Simplifies a fraction to its lowest terms and returns it as a mixed number. |
| modular_inverse_sum | Calculates the sum of modular inverses of the given expressions modulo the specified modulus. |
| simplify_mixed_numbers | Simplifies the sum of two mixed numbers and returns the result as a string in the format ’a b/c’. |
| sum_of_digit_factorials | Calculates the sum of the factorial of each digit in a number. |
| sum_of_primes_below | Calculates the sum of all prime numbers below a given threshold. |